Please assist me in helping to solve the following problem:
3x + 2y = 3
4x -3y = -13
Your analysis and comments are appreciated. Thank you.

Answer by stanbon(57377)

solve the following problem:
3x + 2y = 3
4x -3y = -13
--------------------
If you want to solve by elimination you need to
have the same number of some variable in the
numerator and in the denominator.
You can pick either of the two variables but
you must create the same number in both equations.
---
Lets say you pick "x".
-----
Multiply thru the 1st equation by 4
Multiply thru the 2nd equation by 3
-----
12x + 8y = 4*3
12x - 9y = 3*-13
----------------------
Subtract the 2nd equation from the 1st to get:
17y = 12--39
17y = 51
Divide both sides by 16 to get:
y = 3
---
Substitute that into either of the original equations
and solve for "x":
3x + 2y = 3
3x + 2*3 = 3
3x + 6 = 3
3x = -3
x = -1
-------
Check the solution in the "other" original equation:
4x -3y = -13
4(-1)-3(3) = -13
-4-9 = -13
-13 = -13
===============
Solution:
x = -1 ; y = 3
=======================
Cheers,
Stan H.